OpenStudy (anonymous):

Pls help!! I need to get a 90 on my precal final to get an A, and this is pretty much the only problem i dont know how to do. State whether the given measurements determine zero, one, or two triangles. C = 37°, a = 16, c = 14

OpenStudy (anonymous):

bump

OpenStudy (jdoe0001):

bump? hhhee

OpenStudy (jdoe0001):

so... haemm have you covered the law of sines yet?

OpenStudy (anonymous):

yea

OpenStudy (jdoe0001):

I'm thinking on finding the other angles... and see maybe the other side.. to see what they give

OpenStudy (anonymous):

ok, gimme a minute to find that real quick

OpenStudy (jdoe0001):

\(\bf \cfrac{sin(C)}{c}=\cfrac{sin(A)}{a}\implies \cfrac{sin(C)\cdot a}{c}=sin(A)\implies \cfrac{sin(37^o)\cdot 16}{14}=sin(A) \\ \quad \\ sin^{-1}\left[\cfrac{sin(37^o)\cdot 16}{14}\right]=\measuredangle A\)

OpenStudy (jdoe0001):

a bit truncated... so \(\bf \cfrac{sin(C)}{c}=\cfrac{sin(A)}{a}\implies \cfrac{sin(C)\cdot a}{c}=sin(A)\\ \quad \\ \cfrac{sin(37^o)\cdot 16}{14}=sin(A)\implies sin^{-1}\left[\cfrac{sin(37^o)\cdot 16}{14}\right]=\measuredangle A\)

OpenStudy (anonymous):

Yea, so A=~43.455 and B=99.5

OpenStudy (jdoe0001):

so.... that should mean that at least one triangle can be made

OpenStudy (anonymous):

yea for sure, but how do we find if a second one can be made?

OpenStudy (jdoe0001):

hmmm

OpenStudy (jdoe0001):

 meaning that the likelyhood is that when the side of the angle given in this case is smaller than one of the two legs of the other side, you can use another angle 180-"c" to make a 2nd triangle